Calculating partial prescure in a gas mixture A 6.00 L tank at 17.8 °C is filled with 7.95 g of boron trifluoride gas and 5.67g of sulfur hexafluoride gas. You can assume both gases behave as ideal gases under these conditions. Calculate the mole fraction and partial pressure of each gas, and the total pressure in the tank. Be sure your answers have the correct number of significant digits mole fraction: alb boron trituoride partial pressure: mole fraction: partial pressure sulfur hexafluoride 0atrn ? Tokal pressure in tank
Answer #1

Since the gases are ideal, we assume they are independent of each other AND we assume they follow the equation of state PV = nRT

BF3:

n = (7.95 g / 67.82 g/mol) = 0.117 mol

PV = nRT
or, P = nRT / V
or, P = [ (0.117 mol) x (0.08206 L atm mol-1 K-1) x (17.8 + 273.15) K ] / (6.00 L)
        = [ (0.117 mol) x (0.08206 L atm mol-1 K-1) x (290.95 K) ] / (6.00 L)
        = 0.47 atm


SF6:

n = (5.67 g / 146.06 g/mol) = 0.039mol

PV = nRT
or, P = nRT / V
or, P = [ (0.039 mol) x (0.08206 L atm mol-1 K-1) x (17.8 + 273.15) K ] / (6.00 L)
        = [ (0.039 mol) x (0.08206 L atm mol-1 K-1) x (290.95 K) ] / (6.00 L)
        = 0.16 atm

Total pressure in the tank = partial pressure of BF3 + partial pressure of SF6
                                         = 0.47 atm + 0.16 atm
                                         = 0.63 atm


Total moles = 0.117 mol (BF3) + 0.039 mol SF6
                   = 0.156 mol

mole fraction BF3 = 0.117 mol / 0.156 mol = 0.75
mole fraction SF6 = 0.039 mol / 0.156 mol = 0.25